Subatomic Shirts' standard cost included 0.70 metres of direct material per t-shirt at $3.80 per metre. Subatomic actually used 32,500 metres of direct material at a cost of $3.75 per metre to produce 50,000 t-shirts.

a) Calculate the direct materials price variance: $

b) Is the direct materials price variance favourable or unfavourable?

c) Calculate the direct materials quantity variance: $

d) Is the direct materials quantity variance favourable or unfavourable?

e) Calculate the total direct materials variance: $

f) Is the total direct materials variance favourable or unfavourable?
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (3.80-3.75)*32500

= 1625

FAVOURABLE

Materials Quantity Variance = (SQ-AQ) * SP

= (0.70*50,000 - 32,500)* 3.80

= 2500

FAVOURABLE

Total direct materials variance

= 1625 F + 2500 F

= 4125

FAVOURABLE

.

.

Comment if you face any issues